# Learning to Disentangle Latent Reasoning Rules with Language VAEs: A Systematic Study

### v1

@@ 5.1 Decoding Evaluation
Table 3: Quantitative evaluation for VAE model with or without BM25 in Math Reason task. We can observe that selecting the training sample using BM25 can help model performance.
Figure 3: Gradient heatmap for the last posterior encoder layer (query_add setup), where the left: math derivation, middle: explanatory reasoning, right: syllogistic reasoning, Top: cls_weight is 1.0, bottom: cls_weight is 0.1. We can observe that the non-diagonal values are notably close to 0 when providing higher cls_weight (the red colour elements are less scattered), suggesting that incorporating rule information during training enhances the separation of rule subspaces in the encoder’s parameter space. We provide the heatmaps of all layers in Appendix B.
5.2 Encoding Evaluation
@@
As we illustrated in Section 3.2 (Equation 9), by measuring the cosine similarity between gradient vectors associated with different rules, we can quantify the separability between different rule subspaces in the encoder, comparing settings with strong cls_weight = 1.0 (left) and weak cls_weight =0.1 (right) during training. As illustrated in Figure 3, when the classification weight (cls_weight) is set to 1.0, most non-diagonal values are close to zero (orthogonality). In contrast, with cls_weight set to 0.1, a greater number of non-diagonal elements exhibit higher values (the red colour elements are much more scattered). This observation suggests that explicit supervision facilitates the separation of reasoning rules within the encoder’s parameter space (Finding).
@@
In Section 3.2, we illustrated that rule separation in the parametric space leads to corresponding separation in the feature space. As shown in Figure 4, the sentence representations tend to form distinct clusters that reflect rule information when the classifier is given a higher cls_weight. However, when the cls_weight approaches zero, these rule-based clusters disappear. This suggests that the neural network relies more on the memorisation of lexical combination than on rule-based learning (Finding).
Figure 4: PCA visualisation for query_add injection setup, where top: cls_weight is 1.0, bottom: cls_weight is 0.1. We can observe that the model struggle to learn the rules when the weight is close to zero, indicating the neural network try to deliver reason behaviour via memorisation, rather than rule-based learning. For other injection setups, their visualisations are provided in Figure 12, 13, and 14.
@@
5.3 Case Study for Math Reasoning
@@
First, we analyse how varying the number of training samples for each operation affects the reasoning capabilities of the decoder-only LM (Qwen2.5-0.5B). As illustrated in Figure 5 (left bar plot), it can be observed that when the number of samples per category exceeds 2,000, there is a noticeable decline in accuracy. This suggests that increasing the sample size may introduce greater variability or complexity, potentially disrupting the consistency of each operation. The observation also highlights a limitation of current autoregressive LMs: rather than engaging in rule-based reasoning, they tend to rely on retrieving memorised training instances embedded in their parameters Zhong et al. (2025). Explicitly injecting latent reasoning representations can help mitigate this issue (Finding).
@@
Second, we assess the parametric separation of rule-based information across different model components, specifically the attention (attn) and feed-forward network (ffn) layers, using the same methodology outlined previously. We report the sum of the average diagonal values across all layers. As illustrated in Figure 5 (right bar plot), the ffn layers (right) exhibit a greater tendency to encode rule-based information compared to the attention layers, indicating a more prominent role in capturing structured reasoning patterns (Finding).
6 Conclusion and Future Work
This study serves as a foundational step in exploring the rule-based representation learning under the language VAE architecture for NLI tasks.
We propose a complete pipeline for learning reasoning rules within Transformer-based language VAEs. This pipeline encompasses three rule-based reasoning tasks, a supporting theoretical framework, and a practical end-to-end architecture. The experiment illustrates the following findings: Disentangled reasoning: Under explicit signal supervision, reasoning rules—viewed as functional mappings—can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space. Prior knowledge injection: injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models. Performance bottleneck: In mathematical reasoning tasks using Qwen2.5 (0.5B), increasing sample count doesn’t improve performance beyond a point. Moreover, ffn layers are better than attention layers at preserving the separation of reasoning rules in the model’s parameters. To investigate compositional behaviour beyond memorisation in NLI, in the future, we will focus on current diffusion-based language models, such as Mmada Yang et al. (2025) and Gemini diffusion, which offer greater flexibility for constraining and optimising the decoding process (van Krieken et al., 2025).
7 Limitations
Large scale decoder.
Since our primary motivation is to investigate latent rule separation, we focus on small-scale LM decoders. This choice allows for accessible evaluation of training dynamics, memory capacity, and model updates within the Transformer architecture. Importantly, because the decoder is fully re-trained, the resulting latent space geometry and separation are not constrained by the size or pre-training of the model, enabling a more controlled and interpretable analysis of the learned representations. However, how we can seamlessly integrate the prior rule information into the frozen LLM decoders should be explored in the future.
Composition and Generalisation.
Encoding rule information in the latent space can support compositional reasoning and enhance generalisation in Natural Language Inference (NLI) tasks. However, in this work, we primarily focus on memorisation rather than generalisation. Specifically, we define a rule as a memorised input-output pattern encoded within the latent space. While this approach facilitates controlled reasoning over known patterns, the broader generalisation capacity remains constrained by the architectural limitations of auto-regressive LM decoders. Therefore, we incorporate such prior rule information into the Query component of the attention mechanism in the LM decoder, enabling more effective retrieval through attention over the Key and Value representations. Further investigation is needed to understand how rule-based composition scales to more complex and diverse inference scenarios.
@@ Appendix A Experimental Details
Table 4 describes the reasoning tasks in the experiment.
@@
Max epoch: 10, lr: 3e-5, GPU device: A6000 $\times$ 2.
@@ Appendix B Additional Results
We provide comprehensive heatmaps covering all encoder layers across the full set of tasks, including Math Reasoning: Figure 6 and 7; Explanatory Reasoning: Figure 8 and 9; Syllogistic Reasoning: Figure 10 and 11.
@@
We provide the latent space visualisation across all reasoning tasks in Figure 12, 13, and 14.

### v2

@@ Introduction
Table 3: Quantitative evaluation for VAE model with or without BM25 in Math Reasoning task.
Figure 2: Gradient heatmap for the last posterior encoder layer (query_add setup), where the left: math derivation, middle: explanatory reasoning, right: syllogistic reasoning, Top: cls_weight is 1.0, bottom: cls_weight is 0.1. We can observe that the non-diagonal values are notably close to 0 when providing higher cls_weight (the red colour elements are less scattered), suggesting that incorporating rule information during training enhances the separation of rule subspaces in the encoder’s parameter space. We provide the heatmaps of all layers in the supplementary material.
Figure 3: PCA visualisation for query_add injection setup, where left three: cls_weight is 1.0, right three: cls_weight is 0.1. We can observe that the model struggle to learn the rules when the weight is close to zero, indicating the neural network try to deliver reason behaviour via memorisation, rather than rule-based learning. For other injection setups, their visualisations are provided in Figure 11, 12, and 13 in the supplementary material.
Encoding Evaluation
@@
As illustrated in Section Rule-based Neural Tangent Kernel (Equation 8), by measuring the cosine similarity between gradient vectors associated with different rules, we can quantify the separability between different rule subspaces in the encoder, comparing settings with strong cls_weight = 1.0 (left) and weak cls_weight =0.1 (right) during training. As illustrated in Figure 2, when the classification weight (cls_weight) is set to 1.0, most non-diagonal values are close to zero (orthogonality). In contrast, with cls_weight set to 0.1, a greater number of non-diagonal elements exhibit higher values (the red colour elements are much more scattered). This observation suggests that explicit supervision facilitates the separation of reasoning rules within the encoder’s parameter space (Finding 4).
@@
Since rule separation in the parametric space leads to corresponding separation in the feature space, as shown in Figure 3, the sentence representations tend to form distinct clusters that reflect rule information when the classifier is given a higher cls_weight. However, when the cls_weight approaches zero, these rule-based clusters disappear. This suggests that the neural network relies more on the memorisation of lexical combination than on rule-based learning (Finding 5).
@@
Case Study for Math Reasoning
@@
First, we analyse how varying the number of training samples for each operation affects the reasoning capabilities of the decoder-only LM (Qwen2.5-0.5B). In Figure 4 (left bar plot), it can be observed that when the number of samples per category exceeds 2,000, there is a noticeable decline in accuracy. This suggests that increasing the sample size may introduce greater variability or complexity, potentially disrupting the consistency of each operation. The observation also highlights a limitation of current autoregressive LMs: rather than engaging in rule-based reasoning, they tend to rely on retrieving memorised training instances embedded in their parameters (Zhong et al. 2025). Explicitly injecting latent reasoning representations can help mitigate this issue (Finding 6).
@@
Second, we assess the parametric separation of rule-based information across different model components, specifically the attention (attn) and feed-forward network (FFN) layers, using the same methodology outlined previously. We report the sum of the average diagonal values across all layers. As illustrated in Figure 4 (right bar plot), the FFN layers (right) exhibit a greater tendency to encode rule-based information compared to the attention layers, indicating a more prominent role in capturing structured reasoning patterns (Finding 7).
Conclusion and Future Work
This work serves as a foundational step in exploring rule-based representation learning under the language VAE architecture for NLI tasks. We propose a complete pipeline for learning reasoning rules within Transformer-based language VAEs. This pipeline encompasses three rule-based reasoning tasks, a supporting theoretical framework, and a practical end-to-end architecture. The experiment illustrates the following findings: Disentangled reasoning: Under explicit signal supervision, reasoning rules can be disentangled within the encoder’s parametric space, resulting in distinct clustering of rules in the output feature space. Prior knowledge injection: injecting reasoning information into the Query enables the model to more effectively retrieve the stored Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models. In addition, we found that FFN layers are better than attention layers at preserving the separation of reasoning rules in the model’s parameters. In the future, we will focus on the investigation of diffusion-based language models, such as Mmada (Yang et al. 2025), which can improve flexibility for constraining the decoding process.
@@ Appendix A Experimental Details
Table 4 describes the reasoning tasks in the experiment. For each corpus, the dataset is split into train $60\%$, valid $20\%$, and test $20\%$ sets.
@@
Max epoch: 10, lr: 3e-5, GPU device: A6000 $\times$ 2. The latent dimension is 32 during the experiment following the same setup as Optimus. All other experimental details are provided in the codebase.
@@ Appendix B Additional Results
We provide comprehensive heatmaps covering all encoder layers across the full set of tasks, including Math Reasoning: Figure 5 and 6; Explanatory Reasoning: Figure 7 and 8; Syllogistic Reasoning: Figure 9 and 10.
@@
We provide the latent space visualisation across all reasoning tasks in Figure 11, 12, and 13.